OpenStudy (anonymous):

One week, Huang ordered 1 pizza and 2 sodas for $12. The next week, he ordered 3 pizzas and 4 sodas from the same restaurant for $30. Write and graph a system of equations to determine the cost of each pizza and the cost of each soda from this restaurant. Let x = the cost of a pizza Let y = the cost of a soda

OpenStudy (anonymous):

x = $ 6 y = $ 3

OpenStudy (anonymous):

x+2y=12 3x+4y=30 Put them in slope intercept form: 2y=-x+12 y=-1/2x+6 4y=-3x+30 y=-3/4x+15/2 Now graph both and you will find that they intersect at (6,3) This means $6 for pizza and $3 for soda.

OpenStudy (anonymous):

System of equations: 1x + 2y = 12 3x + 4y = 30 Solving for x x = -2y + 12 Substituting, 3(-2y + 12) + 4y = 30 -6y + 36 + 4y = 30 -2y = -6 y = 3 Substituting again, x = -2(3) + 12 x = -6 +12 x = 6 Answer x = $6, y = $3